Calculate the power of a pump which lifts 500kg of water through a vertical height of 400cm in 5sec (assume g to be 10 m/s).
1
Expert's answer
2019-07-29T08:24:33-0400

Work must be done

"W=mgh"

Power it is

"P=\frac{W}{t}"

So, The power of pump

"P=\frac{mgh}{t}"

Calculating

"P=\frac{500\cdot 4\cdot 10}{5}=4 (kW)"